Exercise 2-3 (Essay) Will Jones, LLP is a small CPA firm that focuses primarily on preparing tax returns for small businesses. The company pays a $500 annual fee plus $10 per tax return for a license to use Mega Tax software. Assume the cost per return is $11.68 for 300 returns and $11.16 for 432 returns and $10.93 for 536 returns. Why does the cost per return differ at each of the three volume levels?
Answer #1

In this problem, there are two types of costs, a fixed cost of $500 annually and a variable cost of $10 per return

As the volume increases, variable cost per unit remains the same but the fixed cost gets divided among more number of returns. Due to this, fixed cost per unit falls and variable cost per unit remains the same. So, the total cost per unit falls with increase in volume

This can be explained with the following table

Calculations A B C = A x B D E = D / A F = B + E
Volume Variable cost per unit Total variable cost Fixed costs Fixed costs per unit Total cost per unit
300 10 3000 500 1.67 11.67
432 10 4320 500 1.16 11.16
536 10 5360 500 0.93 10.93